OpenStudy (anonymous):

A mother wants to invest $13,000 for her children's education. She invest a portion of the money into a bank certificate of deposit (CD Account) which earns 4% and the remainder in a savings bond which earns 7%. if the total interest earned after one year is $780, how much money was invested at each rate? How much money was invested in the CD account? How much money was invested in the saving bonds?

OpenStudy (anonymous):

A mother wants to invest $13,000 for her children's education. She invest a portion of the money into a bank certificate of deposit (CD Account) which earns 4% and the remainder in a savings bond which earns 7%. if the total interest earned after one year is $780, how much money was invested at each rate? How much money was invested in the CD account? How much money was invested in the saving bonds?

OpenStudy (anonymous):

let the sum invested in CD a/c = x then sum invested in S.Bond = 13000 -x In ONE year interest earned on CD a/c = 0.04*x Interest earned on S.Bond = 0.07(13000-x) = 910 - 0.07x ATQ, 0.04x + (910 - 0.07x) = 780 => - 0.03x + 910 = 780 => - 0.03x = 780 - 910 => x = - 130/-0.03 => x = 4333.33 so sum invested in CD a/c = $ 4333.33 sum invested in S.Bond = $ 8666.67 interest earned on CD a/c = $ 173.33 interest earned on S.Bond = $ 606.67